OpenStudy (anonymous):

Logs!! I need to know the followingg about F(x) = 62.13e^(-0.55x) •The logarithmic function based on the graphed function. •A description of what the logarithmic function would look like. •The half-life function Thanks!

OpenStudy (anonymous):

is the first question asking "what is the inverse of this function?"

OpenStudy (anonymous):

No, I graphed points and then I drew a function by connecting them and now I need to know what it would be in terms of log.

OpenStudy (anonymous):

then i am afraid i do not understand the question because you have an exponential function, not a logarithmic one we can find the logarithmic function that is the inverse of this one

OpenStudy (anonymous):

Oh, then maybe that is what I need to do, I guess I am supposed to put it intosomething along the lines of In y = -0.55x + 4.13

OpenStudy (anonymous):

ok we can do that

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[F(x) = 62.13e^{-0.55x} \] put \[x=62.13e^{-0.55y}\]and solve for \(y\) divide by 62.13 to get \[\frac{x}{62.13}=e^{-0.55y}\] take the log to get \[-.55y=\ln(\frac{x}{62.13})\] divide by -.55 now here \(-.55=-\frac{55}{100}=-\frac{11}{20}\) so dividing by -.55 is the same as multiplying by \(-\frac{20}{11}\) so we get \[y=-\frac{20\ln(\frac{x}{62.13})}{11}\]

OpenStudy (anonymous):

Thanks!! Also, I need help with the description of the log, Thanks so much!

OpenStudy (anonymous):

What it would look like on a graph

OpenStudy (anonymous):

if you have the graph of \(F\) it would be symmetric with respect to the line \(y=x\)

OpenStudy (anonymous):

Awesome! Thankss!! :)

OpenStudy (anonymous):

i should say "rotated about the line \(y=x\)" yw